OpenStudy (barrelracing):

Is the expression x3•x3•x3 equivalent to x3•3•3? Why or why not? Explain your reasoning.

OpenStudy (alekos):

\[x^{a}.x^{b} = x ^{a+b}\]

OpenStudy (barrelracing):

?

OpenStudy (barrelracing):

help please

OpenStudy (barrelracing):

@Michele_Laino

OpenStudy (alekos):

do you mean x^3.x^3.x^3 ?

OpenStudy (barrelracing):

yes

OpenStudy (michele_laino):

hint: \[\Large {x^3} \cdot {x^3} \cdot {x^3} = {x^{3 + 3 + 3}} = ...?\]

OpenStudy (barrelracing):

I don't understand

OpenStudy (michele_laino):

it is the application of the property of multiplication of powers with the same basis, please see the post of @alekos above

OpenStudy (barrelracing):

I just don't exactly know what im suppose to do.

OpenStudy (michele_laino):

I think that you have to compute the multiplication of those powers, namely x^3 * x^3 * x^3

OpenStudy (barrelracing):

x^27

OpenStudy (michele_laino):

no, since we have this: \[\Large {x^{27}} = {x^{3 \times 3 \times 3}} = {\left\{ {{{\left( {{x^3}} \right)}^3}} \right\}^3}\]

OpenStudy (barrelracing):

so they are equivalent?

OpenStudy (michele_laino):

no, they are different, since this quantity: \[\Large {x^3} \cdot {x^3} \cdot {x^3} = {x^{3 + 3 + 3}}\] is the multiplication of powers whereas this quantity: \[\Large {x^{3 \times 3 \times 3}} = {\left\{ {{{\left( {{x^3}} \right)}^3}} \right\}^3}\] is the power of power of power

OpenStudy (michele_laino):

and when x is different from 1, those quantities are different

OpenStudy (barrelracing):

ok thank you

OpenStudy (michele_laino):

thanks! :)